You work at a T-shirt printing business. 564 of 4,000 T-shirts are printed improperly. Here is the question:
Predict the number of improperly printed T-shirts in a batch of 2,000.


Any help would be great!-Jasmine

Answer by checkley71(8403)

564/4000=x/2000 cross multiply
4000x=564*2000
x=(564*2000)/4000
x=564/2
x=282 improperly printed out of a pathch of 2,000 shirts.